OpenStudy (anonymous):

What is the value of x in the equation.... (logarithms), equation will be posted in comments

OpenStudy (anonymous):

\[\log_{8}x + \log_{4}x = 10/3 \]

OpenStudy (kc_kennylau):

Firstly, the law: \[\Large\log_ab=\dfrac{\log_nb}{\log_na}\] Substitute \(\Large a=8\), \(\Large b=x\) to get: \[\Large\log_8x=\dfrac{\log_nx}{\log_n8}\]

OpenStudy (kc_kennylau):

Substitute \(\Large a = 4\), \(\Large b = x\) to get: \[\Large\log_4x=\dfrac{\log_nx}{\log_n4}\]

OpenStudy (kc_kennylau):

What would be a reasonable \(\Large n\) to fit in?

OpenStudy (anonymous):

10?

OpenStudy (kc_kennylau):

\[\Large \log_8x+\log_4x=\dfrac{10}3\] \[\Large \dfrac{\log_nx}{\log_n8}+\dfrac{\log_nx}{\log_n4}=\dfrac{10}3\] I would actually choose \(\Large n = 2\) because \(\Large 4=2^2\) and \(\Large 8=2^3\)

OpenStudy (kc_kennylau):

Let's use \(\Large n = 10\) anyway.

OpenStudy (kc_kennylau):

\[\Large \dfrac{\log_{10}x}{\log_{10}2^3}+\dfrac{\log_{10}x}{\log_{10}2^2}=\dfrac{10}3\]

OpenStudy (anonymous):

I was debating whether to use 2, but I didn't know how to use the 2 to the power of 2 and 3 in the equation.

OpenStudy (kc_kennylau):

okay, let's use \(\Large n = 2\) then.

OpenStudy (kc_kennylau):

\[\Large \dfrac{\log_2x}{\log_22^3}+\dfrac{\log_2x}{\log_22^2}=\dfrac{10}3\]

OpenStudy (kc_kennylau):

\[\Large \dfrac{\log_2x}3+\dfrac{\log_2x}2=\dfrac{10}3\]

OpenStudy (kc_kennylau):

Are you able to find \(\Large \log_2x\) now?

OpenStudy (anonymous):

Sorry, I'm not completely sure how to finish it.

OpenStudy (anonymous):

Would we simplify them into one fraction?

OpenStudy (kc_kennylau):

Let's substitute \(\Large a \) as \(\Large \log_2x\).

OpenStudy (kc_kennylau):

The equation then becomes: \[\Large\frac a3+\frac a2=\frac{10}3\]

OpenStudy (anonymous):

a=4?

OpenStudy (kc_kennylau):

Yes :)

OpenStudy (anonymous):

x=16!

OpenStudy (kc_kennylau):

So \(\Large \log_2x=4\)

OpenStudy (kc_kennylau):

Yes :D

OpenStudy (kc_kennylau):

Smart boy/girl

OpenStudy (anonymous):

Haha, girl - Thank you so much!

OpenStudy (kc_kennylau):

no problem :pp